# Using Half-Precision for GNN Training

### v1

@@ 4.2. SDDMM Reduction: More Vector Data-Types
SDDMM reduction happens across feature dimension that requires inter-thread communication for reducing the vector dot product to a single scalar. Inter-thread communication implies a memory barrier, and hence it does not allow the compiler to reorder memory instructions (e.g., data loads) across the memory barrier, and makes warp threads wait here until all threads finish their instruction issue. Hence, threads only issue one data-load if using half or half2, each thread loads only 2 half features. Further, a warp performs four rounds of inter-thread communication to reduce 32 features when using half2 data-type.
@@
Thus the baseline version of SDDMM using half2 data type does not have an optimal design. Reliance on half2 alone is not sufficient to realize the full data-load potential of GPUs as we show in the evaluation that half2 cannot achieve the desired performance. HalfGNN proposes new optimizations to fully take advantage of half-precision in terms of two new vector data types, as discussed next.
@@
Half4 and Half8. We propose, for the first time, two new vector data types, half4, and half8 similar to float2 and float4 respectively. The proposed half4 provides parity with float2 where 4 half-precision features are grouped as one half4 data point. Similarly, half8 treats 8 consecutive half-precision features as one half8 feature. However, there isn’t native support for these two data types, though CUDA supports float2 and float4 respectively. Hence, the proposed half4 internally uses float2 data-type to issue data-load, i.e. each warp can issue 256 bytes of data-load in a single data-load instruction (4×4$\times$4 × the number of features compared to float). Similarly, half8 internally relies on float4 data type to issue 512 bytes of data-load using a single data-load instruction (8×8$\times$8 × the number of features to float). Continuous memory locations of features can be accessed using a pointer data types of half, half2, half4, and half8 for read-write operation memory constraints are maintained. For example, if we have a pointer pointing to an array of data-type half, we can easily switch its pointer type from half to half2, half4, and half8 if the array size is a multiple of 2, 4, and 8 respectively. Thus, HalfGNN offers a range of vector data types to researchers.
For SDDMM, we use half8 data-type wherever possible. It then performs a dot product of vertex features of rows and column IDs of each NZE. Using half8, each thread loads 8 half features of columns, minimizes the inter-thread communication as half8 data-type requires only two rounds of inter-thread communication as only 4 threads are needed to handle 32 features compared to 4 rounds in half2-based SDDMM. Hence, by using half8 (i.e. 4 half2) for data-load, we observe an increase in instruction-level parallelism (ILP) compared to using half2 for feature load as each thread can now issue 4 half2 feature load before waiting at the memory barrier. Better ILP leads to better performance. The sub-warp concept( section 4.1.1) helps us utilize the idle threads better when using these vector types.
@@
However, please note that, like the float2 and float4, half4 and half8 do not improve computation throughput than half2. Hence, even though we provide arithmetic on half4 and half8 data types, it internally relies on half2 arithmetic.
4.3. SpMM Reduction
For SpMM, one can also use half4 or half8, though it does not suffer from the issues that we outlined for SDDMM due to the absence of inter-thread communication. Specifically, reduction in SpMM is performed in the neighborhood dimension, so we need inter-thread communication only when the whole row is not assigned to a single warp. Indeed, Float-based SpMM has achieved workload balancing by dividing the whole neighborhood of any row to multiple warps. Fig. 3 shows the proposed edge-parallel HalfGNN SpMM, which we explain shortly. However, we observe that almost all state-of-the-art workload-balanced SpMM follow these rules:
@@ 4.3. SpMM Reduction
    More than one edges are allocated to the same warp or CTA, whereby threads handle them in a loop. The consecutive edges follow the spatial ordering hence their row ID may be the same or in monotonically increasing order.
@@
Rule 2 implies that a warp or CTA handles only a group of edges at any time, and is the main distinguishing factor between vertex-parallel and edge-parallel workload-balanced solutions. Specifically, in an edge-parallel design, a row-split may be observed within a warp for the latter as shown in Fig. 3. In a vertex-parallel design, rowID remains the same. For example, Huang et al (huang2021understanding, ) and GNNAdvisor (wang2021gnnadvisor, ) form a batch of 32 neighbors operated by each warp. Even vanilla vertex-parallel SpMM that offers no workload balancing, such as GE-SpMM (gespmmsc2020, ) has an implicit batching where each warp handles 32 neighbors of any vertex before processing the next 32 batches of neighbors for the same vertex in the next iteration, except the last iteration where it may process less than 32 neighbors depending on the vertex degree.
Irrespective, the design implies that workload-balanced solutions are easily amenable for discretized reduction scaling to protect against overflow, as discussed next.
4.3.1. Discretized Reduction Scaling
The proposed method is based on the observation on how the reduction is performed currently, and how can it be tailored to solve the problem. In effect, SpMM + degree-norm is treated as a custom SpMM kernel with ‘mean’ as a reduction operation, thus solving the problem for both GCN and GIN models.
We term the currently used reduction approaches as post-reduction scaling (degree-norm scaling applied once at the end of the reduction of the full neighborhood of vertices), performing exactly one division per vertex at the end of reduction. The other end of the spectrum would be to perform degree norm scaling before reduction, i.e., divide each dot product by degree and then perform reduction, pre-reduction scaling. This would perform exactly $|N(v)|$ count of division for each vertex, would avoid INF generation, but would introduce more arithmetic operations. While a few values might vanish (underflow) due to division at each dot product.
The proposed discretized reduction scaling shares the pros of both and the cons of none, as degree-norm scaling applies after reduction on a batch of neighbors for each vertex (rule 2). To illustrate, we perform the degree-norm scaling at the end of the reduction for each batch. Thus discretizing the reduction never risks getting INF as long as the the maximum number of neighbors in the batch is small. As one iteration of computation for a row handles a handful of neighbors, each iteration does produce a partially reduced value that never overflows on which the degree-norm scaling is performed. Thus, the proposed design changes are minimal and effective.
@@ 4.3.1. Discretized Reduction Scaling
Additional Overflow in GIN. Discretized Scaled SpMM completely removes accuracy issues for GCN, but it is still not enough for GIN. In GIN aggregation, we observe 2 parts, one is the right representing unweighted SpMM being added to scaled source node features as shown in Eq. 3. The overflow issue does occur when adding self features for sources to SpMM output. We introduce a non-learnable parameter $\lambda$, to scale the aggregation down. Ideally, $\lambda=0.1$ worked fine for all our robust testing. So, we propose the following kind of aggregation for GIN in half-precision:
@@ 4.3.2. Non-Atomic Writes
The discretized reduction also introduces conflicting write for which prior float-based works have relied on atomic write for the data-store phase, which is very costly for half2-based atomic writes. We first lay out when conflicting writes occur and how to resolve them: a) if one or more rows are contained exclusively within a warp, there is never any conflicting write, and hence no atomic write is needed. b) If one or more rows are contained exclusively within a CTA, then an intra-CTA communication can resolve those conflicting writes locally within the CTA to generate non-conflicting writes. We discuss this soon. c) Only when a row extends to more than one CTA that conflicting writes cannot be resolved. To resolve this case, we work with point (b) above and deploy a staging buffer to direct the conflicting write. The staging buffer is then re-read in a separate follow-up kernel to write to the output tensor. Fig. 3 shows the details, discussion follows next.
Conflict Write Handling. The thread-local reduced value is written non-atomically to output tensor directly in case of row-split as shown for the yellow row in Fig. 3. The reduced values at warp-end (non-row-split case) are the ones that are resolved first locally within a CTA by using shared memory. Kindly note that current libraries, such as Nvidia’s CUB offer intra-CTA communication only for scalar dot products for reduction, and are insufficient for exchanging vector dot products.
@@
Figure 4. HalfGNN achieves the same accuracy as float-based DGL, confirming the applicability of the proposed techniques. Only labeled datasets are used to show accuracy.
@@
We propose a new design for intra-CTA communication, which also handles the case of sub-warp (section 4.1.1). In such a case, not all threads need to communicate with every other thread; the communication happens between a thread of a sub-warp to the same thread ID of all other sub-warps of the CTA so that the same feature gets reduced. Hence, if there are 8 sub-warps within a CTA, we only perform three rounds of communications within the CTA. The proposed library relies on cached row ID stored in the shared memory to know if the dot products stored in a sub-warp should be merged to the dot products of another sub-warp.
The resolved writes are written directly to the output tensor. The unresolved writes are the ones that can only be resolved by looking into the writes by the next CTAs. We treat the last (partial) row assigned to this CTA as the conflicting write. This generates just one final conflicting write of size feature-length, which is the output corresponding to the row ID of the last edge(s) assigned to this CTA and are written to the staging buffer. A follow-up kernel is written to handle this buffer to perform the final write.
Staging Buffer & Writing Heuristic. This buffer is where we store the carry-outs is of size $|CTA|\times 4$ bytes. The number of total destination vertices is $|CTA|$ and the memory requirement is $|CTA|\times 4$ bytes. This combined data structure requires little memory, and we save a lot by using half-precision for state tensors already.
4.3.3. Generality of the Solution
Our non-atomic design principle involves intra-CTA, & inter-CTA conflict resolution, with 2-stage kernel design. This applies to edge-parallel design as HalfGNN itself is edge-parallel. Not only that, this applies to even vertex-parallel designs. We apply the vectorization design and non-atomic write to Huang et al  (huang2021understanding, ) and gain 1.79×1.79$\times$1.79 × speedup on average for SpMM. This shows the proposed optimizations can massively improve prior works.
@@ 4.4. Shadow APIs
In mixed-precision training, DL frameworks invoke float versions of many operations (PyTorchAMP, ), e.g. exponential (section 3.1.2). However, if we are sure that such operations would produce the output within the numerical range of half-precision, we advocate using the half-precision version. This guarantees an optimal number of kernel calls are half-precision and the least number of float conversions for maximum performance.
However, DL frameworks, such as PyTorch, do not automatically invoke such operation in half-precision (PyTorchAMP, ). To this end, we provide similar APIs, called shadow API that guarantee that the half-precision version would be invoked if the input tensor is of half data-type, and the float version would be invoked if the input tensor is of float type. Hence, one can always use the proposed shadow API if we are sure that overflow would not incur in half-precision.
5. Evaluations
The following experiments aim to comprehensively evaluate the proposed optimizations and perform a comparative study compared to prior GNN works. All the experiments are conducted in NVIDIA A100 GPU which has 40GB of GPU memory. We conducted our experiments on 16 datasets mentioned in Table 1. Only 5 datasets (Cora, Pubmed, Citeseer, Ogb-product, and Reddit) are labeled where accuracy is measured, whose input feature lengths are shown. Others are used for evaluating runtime, and memory usage only, hence they contain generated features.
Table 1. Graph datasets. * denotes labeled dataset. F = Input feature-length, C = Prediction categories

### v2

@@ 5.1.1. Observation
Inter-thread communication implies a memory barrier that does not allow the compiler to reorder memory instructions (e.g., data loads) across the barrier. The barrier makes warp threads wait until all threads finish their work. Hence, threads only issue one data load of size 32 bits (half2) before synchronizing at the barrier. Further, a warp performs four rounds of inter-thread communication to perform reduction on 32 half2 values.
@@
Thus, the baseline version of SDDMM using half2 data type does not have an optimal design as it is not sufficient to realize the full data-load potential of GPUs, as we show in the evaluation (section 6.3). HalfGNN proposes two new vector data types to fully optimize the performance, as discussed next.
5.1.2. Half4 and Half8
We propose, for the first time, two new vector data types of half4 and half8 that pack four and eight half-precision data. To provide their native support, half4 internally uses float2 data-type to issue data-load, i.e., each warp can issue 256 bytes of data-load in a single data-load instruction where each warp thread loads 4 half-precision data. Similarly, half8 internally relies on float4 data type to issue 512 bytes of data-load using a single data-load instruction. Float2 and float4 are natively supported by GPUs for data-load purposes and not for computation. Thus, half4 and half8 use half2 for arithmetic.
@@ 5.1.2. Half4 and Half8
To this end, HalfGNN offers a range of data types to researchers where continuous memory locations of features can be accessed, using a pointer data type of half, half2, half4, and half8 for read-write operation, provided memory constraints are maintained. For example, if we have a pointer pointing to an array of data-type half, we can easily switch its pointer type from half to half2, half4, and half8 if the array size is a multiple of 2, 4, and 8 respectively.
5.1.3. Half8 for SDDMM
HalfGNN SDDMM follows an edge-parallel design and continues to have the two-phase data-load design. However, it uses half8 in phase 2 of data-load. Using half8, each warp thread loads 8 half features of the non-zero columns in phase 2 of the data-load (section 4.1.2). It then performs a dot product of the vertex features of rows and column IDs of each NZE. The dot product phase uses half2 because arithmetic on half8 is not supported by hardware. The usage of half8 also minimizes the inter-thread communication. To illustrate, consider 32 as feature-length, where only 4 threads are needed. This requires only two rounds of inter-thread communication. The half2-only solution requires four rounds of inter-thread communication, while a half-based SDDMM would use five rounds of inter-thread communication.
The major reason for the increased performance of half8-based SDDMM is the lesser impact of the memory barrier on the data load. In other words, we now effectively issue a load of 4 half2 features before observing the memory barrier compared to issuing just one half2 feature load in half2-only SDDMM. The sub-warp concept (section 4.1) helps us utilize the idle threads when using these vector types.
@@ 5.1.3. Half8 for SDDMM
Please note that, like the float2 and float4, half4 and half8 do not improve computation throughput compared to half2. Hence, even though we provide arithmetic on half4 and half8 data types, it internally relies on half2 arithmetic. We continue to rely on feature padding to use the newly proposed data type when the feature lengths are not appropriate, such as in the last layer of GNN, where the feature length is determined by the number of classes to predict.
5.2. SpMM: Discretized Reduction Scaling
The reduction in SpMM is performed in the neighborhood dimension, so we need inter-thread communication only when the whole row is assigned to more than one warp– a norm in most workload-balanced SpMM designs (huang2021understanding, ; wang2021gnnadvisor, ; gnnone, ; yang2018design, ; wang2023tc, ). However, inter-thread communication is replaced with atomic write in the prior float-base workload-balanced SpMM solutions, and is a major performance bottleneck. This sub-section shows a close association between this approach and the value overflow issue (section 3.1), to solve both together.
5.2.1. Observation
Fig. 4 shows the edge-parallel HalfGNN SpMM incorporating the proposed discretized reduction scaling, which we explain shortly. The proposal is based on our observation that almost all state-of-the-art workload-balanced SpMM follow these rules:
@@ 5.2.1. Observation
    More than one edges are allocated to the same warp or CTA, whereby threads handle them in a loop. The consecutive edges follow the spatial ordering, hence their row ID may be the same or in monotonically increasing order if a row split is observed.
@@
Rule 2 implies that a warp or CTA handles only a group of edges at any time. We will explain soon how this is done for different computation paradigms of edge-parallel and vertex-parallel SpMM. But first, this count is small enough that we observe that SpMM forms a batching of edges or neighbors. Hence, we claim the computation is discretized, which we exploit to remove atomic writes and protect value overflow.
The key difference between edge-parallel and vertex-parallel is that row-split is observed in the former within a warp, as shown for the third warp of the first CTA in our edge-parallel approach of Fig. 4, as it allocates every warp an equal number of edges to compute irrespective of their source vertex ID (row ID). On the other hand, workload-balanced vertex-parallel SpMM does not force row-split. For example, Huang et al. (huang2021understanding, ) and GNNAdvisor (wang2021gnnadvisor, ) allocate only a group of 32 neighbors to each warp. The last warp may handle fewer than 32 neighbors, as the vertex degree may not be a multiple of 32. Even vanilla vertex-parallel SpMM that offers no workload balancing, such as GE-SpMM (gespmmsc2020, ), has an implicit grouping where each warp handles 32 neighbors of any vertex before processing the next 32 neighbors for the same vertex in the next iteration. The last iteration may process fewer than 32 neighbors, depending on the vertex degree. We, of course, proposed to handle at least 64 edges in place of 32 in such designs (section 4.1.1), nonetheless, they all follow a discretization of computation in place, which we exploit next to protect against value overflow.
5.2.2. Protecting Against Value Overflow
Our proposed method is based on the observation of how the reduction is performed currently and how it can be tailored to solve the problem. In effect, SpMM + degree-norm is treated as a custom SpMM kernel with ‘mean’ as a reduction operation, thus solving the problem for both GCN and GIN models.
We term the currently used reduction approaches as post-reduction scaling (degree-norm scaling applied once at the end of the reduction of the full neighborhood of vertices), performing exactly one division per vertex at the end of the reduction. The other end of the spectrum would be to perform degree norm scaling before reduction, i.e., divide each dot product by degree and then perform reduction, called pre-reduction scaling. This would perform exactly $|N(v)|$ count of division for each vertex, avoiding INF generation, but introducing more arithmetic operations. While a few values might vanish (underflow) due to division at each dot product.
The proposed discretized reduction scaling shares the pros of both and the cons of none. This applies degree-norm scaling after reduction on a batch of neighbors for each vertex (observation rule 2). To illustrate, we perform the degree-norm scaling at the end of the reduction for each batch. Thus, discretizing the reduction never risks getting INF as long as the maximum number of neighbors in the batch is small, which is the case as discussed in the observation. Thus, the proposed design changes are minimal and effective.
@@ 5.2.2. Protecting Against Value Overflow
Additional Overflow in GIN. Discretized Scaled SpMM completely removes accuracy issues for GCN, but it is still not enough for GIN. In GIN aggregation, we observe 2 parts, one is the right representing unweighted SpMM being added to scaled source node features as shown in Eq. 3. The overflow issue does occur when adding self-features for sources to the SpMM output. We introduce a non-learnable parameter $\lambda$, to scale the aggregation down. Ideally, $\lambda=0.1$ worked fine for all our robust testing. So, we propose the following kind of aggregation for GIN in half-precision:
@@ 5.2.3. Non-Atomic Writes
The discretized reduction also introduces conflicting writes for which prior float-based works have relied on atomic writes for the data-store phase, which is very costly for half2-based atomic writes as per our measurements (6.3.2).
Conflict Write Identification and Handling. We now lay out when the conflicting writes occur and how to resolve them in our edge-parallel SpMM (Fig. 4).
  1. (1)
    if one or more rows are contained exclusively within a warp, there is never any conflicting write, and hence no atomic write is needed.
  2. (2)
    If one or more rows are contained exclusively within a CTA, then an intra-CTA communication among its warps can resolve those conflicting writes locally within the CTA to generate non-conflicting writes. We will discuss this soon.
  3. (3)
    Only when a row extends to more than one CTA, the conflicting writes cannot be resolved locally. a) Kindly note that the reduced values generated in the case of row-split within a warp are non-conflicting if we never perform any other conflicting write to the output tensor $Y$ directly. That is, the third warp in Fig. 4 is non-conflicting for row ID $V1$ if the first and the second warps do not perform their write (for $V1$) to $Y$ directly. b) The reduced values at warp-end (non-row-split case) are the ones that are always conflicting.
Fig. 4 shows the details: a) the thread-local reduced value is written non-atomically to the output tensor directly in case of row-split (only) as shown for the yellow row (Case 3(a) above). b) The reduced values at warp-end (non-row-split case), i.e., case 2 and case 3(b) above, are the ones that are resolved first locally within a CTA by using shared memory. This resolves all conflicting writes within CTA, and finally generates one conflicting write of size $|F|$ per CTA: the ones that can only be resolved by looking into the writes by the next CTAs. The former is written directly to $Y$ while the latter is written to the staging buffer. The staging buffer is then re-read in a separate follow-up kernel to write to the output tensor.
@@
Figure 5. HalfGNN achieves the same accuracy as float-based DGL, confirming the applicability of the proposed techniques. Only labeled datasets (Table 1) are used to show accuracy.
@@
Kindly note that current libraries, such as Nvidia’s CUB, offer intra-CTA communication only for scalar dot products for reduction, and are insufficient for exchanging vector dot products, as needed for our purpose. We propose a new design for intra-CTA communication, which also handles the case of sub-warp (section 4.1). In such a case, not all threads need to communicate with every other thread; the communication happens between a thread of a sub-warp to the same thread ID of all other sub-warps of the CTA, so that the same feature gets reduced. Hence, if there are 8 sub-warps within a CTA, we only perform three rounds of communications within the CTA. The proposed library relies on the cached row ID stored in the shared memory to know if the dot products stored in a sub-warp should be merged with the dot products of another sub-warp.
Staging Buffer Size. This buffer stores the conflicting write of each CTA. Hence, its size is $|CTA|\times|F|$ features. This buffer requires little memory and is reused by each SpMM. Further, we already save a lot of memory by using half-precision for state tensors, hence this small buffer allocation is not a major concern.
@@ 5.3. Shadow APIs
In mixed-precision training, DL frameworks invoke float versions of many operations (PyTorchAMP, ), e.g., exponential (section 3.1.2). However, if we are sure that such operations would produce the output within the numerical range of half-precision, we advocate using the half-precision version. This guarantees an optimal number of kernel calls are half-precision and the least number of float conversions for maximum performance.
However, DL frameworks, such as PyTorch, do not automatically invoke such an operation in half-precision (PyTorchAMP, ). To this end, we provide similar APIs, called shadow API, that guarantee that the half-precision version would be invoked if the input tensor is of half data type, and the float version would be invoked if the input tensor is of float type. Hence, one can always invoke the proposed shadow API inside the graph convolution layer if they are sure that overflow would not incur in half-precision.
5.4. Generality of the Solution
Most of the proposed ideas are equally applicable to edge-parallel, vertex-parallel, and their variants, though HalfGNN is an edge-parallel solution. For example, usage of vector data-types (half2, half4, and half8), edge feature mirroring, feature padding, sub-warps, discretized SpMM, removal of atomic writes, and proposed shadow APIs apply to all. The key differences are in phase 1 of data load, where edge features are loaded, and in the handling of conflicting writes, which are discussed next.
Data-Load of Edge features. An edge-parallel system allocates an equal number of edges to every warp. However, vertex-parallel systems allocate unequal edges per warp as they never experience row-split. For example, a few workload-balanced vertex-parallel SpMM (huang2021understanding, ; wang2021gnnadvisor, ) allocates a maximum of 32 neighbors of a vertex to each warp to compute on. However, the last warp allocated to process such vertices may process fewer than 32 neighbors, depending on the degree of the vertex. Similarly, warps that handle vertices with a degree smaller than 32 process fewer neighbors. Though their allocation strategy impacts their workload balancing, it affects the data load of edge features, which does not always start at an even offset. To illustrate, if the first vertex has 11 neighbors, the memory pointer pointing to the edge features corresponding to the next vertex’s neighborhood cannot be type-cast to half2 due to the memory alignment issue, leading to complexity in using half2.
Conflict Write Handling. Our non-atomic design principle involves intra-CTA and inter-CTA conflict resolution, with a 2-stage kernel design. This applies to edge-parallel design as HalfGNN itself is edge-parallel. This applies to vertex-parallel design as well, except that we do not require intra-warp conflict resolution because a row split is not observed in this case. However, intra-CTA and inter-CTA remain in place.
Implementing Half-precision SpMM of Haung et al. (huang2021understanding, ). To tackle the memory alignment issue, we start edge-feature fetch from one position earlier if needed, and handle this carefully during edge-feature mirroring to write the correct value in the shared memory. We apply the rest of the vectorization design and non-atomic write. section 6.3.3 shows the benefit of this design.
Table 1. Graph datasets. * denotes labeled dataset while others have generated features and labels. $|F|$ is the Input feature-length, $|C|$ is the Prediction categories (or output feature-length), while the intermediate feature-length is set as 64.
@@ 5.4. Generality of the Solution
6. Evaluations
We comprehensively evaluate the proposed optimizations and perform a comparative study over prior GNN works. All the experiments are conducted on an NVIDIA A100 GPU (40GB of GPU memory) using 16 datasets mentioned in Table 1. Only 5 datasets (Cora, Pubmed, Citeseer, Ogb-product, and Reddit) are labeled. HalfGNN [^1] Code can be accessed from https://github.com/the-data-lab is integrated into the GNNBench benchmarking platform (GNNBench, ) to avoid pitfalls that have been discovered recently (Graphpy, ). The platform allows to use of non-labeled datasets by using generated labels and features, whose dimensions are listed in Table 1, and are used for only performance measurements due to the limited number of labeled datasets. It also recommends that runtime results on smaller datasets (G0–G2) are not reliable, hence, they are used only for verifying accuracy.
We focus on GCN, GAT, and GIN models. These models contain intermediate feature lengths of 64. The intermediate feature length is important as the sparse kernels usually have this feature length during computation. The original feature length is usually projected to this intermediate feature length through GeMM (a part of the linear layer). Also, the SpMM and SDDMM of the last layer usually use a feature length equal to the prediction categories ($|C|$ in Table 1). Thus, our evaluations include various feature lengths for sparse kernels. All GNNs were trained for 400 epochs.
Kindly note that DGL is the state-of-the-art GNN system for half-precision, though it suffers from value overflow in GCN and GIN models, which directly uses Cusparse half-precision SpMM. Pytorch-Geometric (fey2019fast, ) frequently runs out of memory when using its COO-based design. Its recent implementation that uses CSR format supports GCN-like models only, which also suffers from data conversion issues, while GAT continues to rely on the COO format (Graphpy, ). Hence, we mostly use DGL as the baseline and rely on various micro-benchmarks and performance counters to justify the proposed optimizations. We also used Huang et al. (huang2021understanding, ) SpMM, a workload-balanced solution, which is the state-of-the-art vertex-parallel float-based SpMM, to show the applicability of the proposed techniques to the vertex-parallel computation paradigm.